# Token-Importance Guided Direct Preference Optimization

### v1

@@ Abstract
Ensuring that large language models (LLMs) generate outputs aligned with human preferences is important for safe and effective AI interactions. While Direct Preference Optimization (DPO) employs an implicit reward function to optimize the policy model, however, it and its related variants overlook the differential importance of individual tokens and are sensitive to judgment noise in preference datasets during generation. Although recent methods attempt to assess the important weight of tokens via probability prediction or simplistic weighting schemes, these evaluation methods are prone to biases and still cannot fully address these issues. To solve this problem, we propose the Token-Importance Guided Direct Preference Optimization (TI-DPO), which introduces two key innovations: the gradient-based token-importance weights that dynamically prioritize critical tokens, and a triple loss that explicitly guides model outputs to approach human-preferred responses and stay away from non-preferred responses. Experimental results show that TI-DPO achieves higher accuracy and stronger generative diversity, providing more stable and computationally efficient solutions compared with DPO and other RLHF methods.
@@ 1 Introduction
LLM has shown proficiency in Natural Language Processing (NLP) [1, 2], logical reasoning [3], and code generation [4], emerging as a focal point of recent research. However, as models may generate outputs inconsistent with intended purposes or ethical standards, human preference alignment aims to ensure that LLMs adhere to human values [5, 6], producing beneficial and harmless content. Against this backdrop, Reinforcement Learning from Human Feedback (RLHF) has become a prevailing approach for achieving alignment [7, 8, 9]. It leverages human-annotated preference data to train reward models and fine-tunes LLMs using Reinforcement Learning (RL) methods [10, 11] like Proximal Policy Optimization (PPO) [12]. While effective, RLHF suffers from inherent limitations: its reliance on iterative policy optimization and complex reward modeling leads to training instability and exorbitant computational costs [13].
Addressing these constraints, Direct Preference Optimization (DPO) [14] eliminates the necessity for a reward model. Instead, it employs direct pairwise preference comparisons to optimise model policies. Inspired by DPO’s implicit reward mechanism, a series of preference optimization models have been proposed in recent years, such as ORPO [7], f-DPO [15] and CPO [16]. However, DPO’s loss function does not account for the varying importance of individual tokens during the generation of preferred responses [17], further intensifying the problem of sampling distribution change. Consider a model generating medical advice: a single misweighted token, for example, "safe" versus "risk", could drastically alter the response’s safety. Traditional DPO treats all tokens equally, potentially overlooking such critical distinctions.
Motivated by these challenges, researchers have proposed token-level variants of DPO such as TDPO [17], Logic-RL [3], RTO [18] and SimPO [19], aiming to decompose preference alignment into fine-grained token contributions. Based on these methods, some works have additionally considered the impact of the token importance weights, but they often rely on probabilistic proxies [20] or simplistic weighting schemes [21], which are prone to bias and fail to fully leverage the structural relationships between semantically similar responses. By contrast, we propose Token-Importance Guided Direct Preference Optimization (TI-DPO), using gradient attribution to dynamically compute token-importance weights. Token importance weights act as a ‘spotlight’, focusing optimization on tokens that most influence human judgments.
Furthermore, recent studies [20, 19] have noticed some shortcomings in DPO’s design: it compares merely "good" versus "bad" samples. Concurrently, LLMs generate preference data through various methods [22], such as rejection sampling [23] and employing different prompts for distinct preference levels [22], which spurred us to define different levels of samples. Under such a background, our TI-DPO method constructs a "good-bad" triplet relationship by incorporating the intermediate generated outputs [24]. This triplet structure explicitly guides the intermediate output to approach human preferences and distance from non-preferred responses, achieving fine-grained preference alignment and promoting a continuous gradient of preference learning, rather than relying solely on binary comparisons.
@@
  • Based on DPO, we propose TI-DPO, a token-level approach with important weights. Integrating triplet loss with gradient-based token importance weighting, TI-DPO focuses optimization on critical tokens, mitigating distribution shift and improving alignment precision.
  • Theoretically, we prove TI-DPO achieves a tighter loss bound than DPO (Theorem 4.2), ensuring more stable optimization. This theorem formally provides a new perspective on comprehending the superiority of TI-DPO in terms of alignment accuracy.
  • Experiment results indicate that TI-DPO surpasses existing methods in aligning LLMs with human preferences. In preference comparisons, TI-DPO achieves the highest accuracy on TruthfulQA and IFEval tasks, and we compare TI-DPO with three base instruction models across multiple task dimensions with radar graphs. Additionally, ablation experiments confirm the necessity of its token-importance guidance and triplet loss components.
@@ 2 Related Work
Human Preference Alignment Human preference alignment has emerged as a critical research paradigm in recent years, focusing on enabling model responses to align with human values and preferences. Early advancements mainly focused on RLHF [26, 27] based on PPO [12]. Due to PPO consuming substantial training resources, some recently proposed RL methods aim to reduce overhead by decreasing parameters of PPO [13] or abandoning the critic model [28]. However, these RL methods may suffer from overfitting in optimal responses. To mitigate this issue, Hu et al. [8] introduced the Reinforce++ model, which employs batch-wise standardized rewards to prevent overfitting and enhance the prompt diversity during training. Concurrently, beyond RL approaches, Rafailov et al. introduced DPO [14], which obviates the need for explicit reward modeling through implicit preference learning. This implicit reward mechanism has inspired a wave of subsequent works [29], such as the SimPO algorithm proposed by Meng et al. [19], which utilizes the sequence-averaged log probability as an implicit reward signal to streamline optimization. Notwithstanding these advancements, DPO’s reliance on large-scale human-annotated preference datasets [30] has motivated derivative studies [31, 9] aimed at reducing data requirements. A notable example is RS-DPO [32], which integrates rejection sampling (RS) with DPO to alleviate data scarcity. However, a more fundamental limitation pertains to the binary nature of traditional preference labels. Although the KTO method proposed by Ethayarajh et al. [33] effectively reduces the reliance on paired preference labels in DPO, most current RLHF and related preference optimizations still mainly rely on binary comparisons between "good" and "bad" responses [34, 7]. Such coarse-grained supervision has obvious shortcomings: human preferences often show continuous gradient differences rather than simply "good" and "bad". Against this backdrop, the triplet optimization method we propose can achieve fine-grained preference alignment.
@@
From Sequence-Level to Token-Level Since RLHF and DPO are effective sequence-level training techniques, KL divergence and reward models are computed at the sequence [16] or sentence level [35], overlooking differences in token importance, which leads to a significant decline in diversity [17]. These limitations have spurred researchers’ research into step [36] or token-level [37, 29] alignment mechanisms, seeking to address the granularity mismatch between coarse-grained sequence rewards and fine-grained token contributions [38]. Lai et al. [39] initiated this research path through Step-DPO, which breaks down the reasoning process into a series of sequential decision points, rather than accessing the entire answer comprehensively. In [17], to address the significant decline in model generation diversity, Token-level DPO (TDPO) reanalyzed and optimized the entire alignment process from the token-level perspective. An additional limitation of RLHF and DPO lies in the fact that rewards are only assigned to the final token, with all other tokens receiving no learning rewards [18]. Meanwhile, Xie et al. [3] proposed a correlation between the frequency of certain tokens and model performance, which inspires us to consider reassigning token weights. In a related vein, Liu et al. proposed Token-level Importance Sampling for DPO (TIS-DPO) [20], which estimates token importance weights using prediction probability differences. Nevertheless, this probabilistic weighting scheme is prone to bias when contrastive models produce inconsistent outputs or fail to capture subtle semantic nuances of human preferences. In contrast, our approach employs gradient-based importance estimation [40], directly quantifying each token’s causal contribution to reward prediction. By focusing on actual gradient impacts rather than probabilistic proxies, this method enhances the precision of aligning model outputs with human preferences.
@@ 3.1 Human Preference Alignment
Here $\sigma(\cdot)$ is sigmoid activation function. Reward model evaluates the LLM’s responses by applying the BT model for ranking losses [26]:
@@
Nevertheless, the RLHF approach has been acknowledged to encounter challenges related to implementation complexity and training instability as highlighted in a study [42]. DPO [14] derived the optimal policy $\pi^{*}_{\theta}$ from the reward model:
$\pi^{*}_{\theta}(y|x)=\frac{1}{Z(x)}\pi_{\text{ref}}(y\mid x)\text{exp}(\frac{1}{\beta}r_{\phi}(x,y)),$ (3)
In Eq.3, $\pi^{*}_{\theta}$ represents the optimal policy, with $\pi_{\text{ref}}$ denoting the initial model. The partition function $Z(x)$ serves to normalize the policy’s probability distribution[14]. The parameter $\beta$ regulates the extent of divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$. DPO rearranges this equation to express the reward as $r_{\phi}(x,y)=\beta\log\frac{\pi_{\theta}^{*}(y|x)}{\pi_{\text{ref}}(y|x)}-\log Z(x)$. Let the input prompt be represented as $x=[x_{1},x_{2},\ldots,x_{m}]$ and the first $t-1$ tokens generated by the model be denoted as $y^{<t}=[y^{1},y^{2},\ldots,y^{t-1}]$. Let $T_{\text{w}}$ and $T_{\text{l}}$ denote the number of preferred tokens and less preferred tokens, respectively. The token-level DPO optimization objective is given by
@@ 3.2 Triplet Loss
Here $[z]_{+}$ denotes the rectified linear unit function, ensuring that it is set to zero if negative. The features extracted from the three inputs are represented by the terms $f\left({x}_{i}\right)$, $f\left({x}_{j}\right)$, and $f\left({x}_{k}\right)$.
@@ 4 Methodology
Driven by the challenges of unstable training and distribution shift in traditional RL alignment methods, we propose the TI-DPO framework. Our key innovation lies in integrating the token-importance weight and triplet loss in the TI-DPO framework, attaining a more optimal equilibrium balance between alignment precision and generative diversity.
@@ 4.1 Token-Level MDP for LLM Preference Alignment
To address the challenges of the sequential and auto-regressive nature of text generation, a token-level Markov Decision Process (MDP) is introduced, which incorporates the notion of token significance to improve the alignment of each token selection with human preferences. This concept is defined through a tuple denoted as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho_{0})$. $\mathcal{S}$ and $\mathcal{A}$ are the state space and action space, respectively. $\mathcal{P}$ is a deterministic transition model among tokens. Here $r$ stands reward model associated with each token, and $\rho_{0}$ indicates the initial state distribution. The initial state is $s_{0}=[x]$, which is simply the input prompt. At each step $t$ of the generation process, the state $s_{t}=[x,y^{<t}]\in\mathcal{S}$ consists of input prompt $x=[x_{1},x_{2},\ldots,x_{m}]$, where $m$ is the count of token, and $t-1$ generated tokens $y^{<t}=[y^{1},y^{2},\ldots,y^{t-1}]$. At each time step $t$, the action $a_{t}=y^{t}$ corresponds to the selection of subsequent tokens for generation. The initial state distribution $\rho_{0}$ represents the distribution over input prompts $x$.
4.2 Gradient Attribution for Token Importance
@@
Building on the token-level MDP framework, we now formalize the calculation of importance weights $w_{t}$. Inspired by the attribution-based rationale extraction from Ballout et al. [40], our approach quantifies token importance through gradient sensitivity analysis, ensuring that critical tokens in human-preferred responses drive the policy optimization process.
Assume that each token $x_{i}$ is transformed into an embedding vector $e_{i}$, forming an embedding sequence $E=[e_{1},e_{2},\cdots,e_{m}]$. We compute the gradient of reward model’s output concerning $e_{i}$ to measure how sensitive the reward is to changes in the token. For each reward $r_{\phi}(s_{j},a_{j})$, the gradient corresponding to $e_{i}$ is $\nabla_{e_{i}}r_{\phi}(s_{j},a_{j})$. Similar to previous work [40], the average gradient for embedding $e_{i}$ is given by:
$\bar{G}_{i}=\frac{1}{t}\sum_{j=1}^{t}\nabla_{e_{i}}r_{\phi}(s_{j},a_{j}).$ (6)
Since $\pi_{\text{ref}}$ is fixed, there is $\nabla_{e_{i}}\log\frac{\pi_{\theta}}{\pi_{\text{ref}}}=\nabla_{e_{i}}\log\pi_{\theta}(y^{t}|x,y^{<t})$. For each token $x_{i}$ in the input or generated sequence, we rewrite (6) with policy log-probability ratio as:
$\bar{G}_{i}=\frac{1}{t}\sum_{j=1}^{t}\nabla_{e_{i}}\log\pi_{\theta}(y^{j}|x,y^{<j}).$ (7)
Although the average operation may weaken the long-range dependence, the generation of each token $y^{t}$ only only depends on the prefix $y^{<t}$, so the gradient $\nabla_{e_{i}}\log\pi_{\theta}(y^{j}|x,y^{<j})$ has implicitly included the historical dependence. We take the average of all time steps to balance the local and global importance. Denote the importance-score for the $i$-th token as $I_{i}$. Following the attribution method of Ballout et al. [40], we calculate $I_{i}$ using the $L_{1}$ norm of gradient:
$I_{i}=\left\|\bar{G}_{i}\right\|_{1}=\sum_{k}\left|\bar{G}_{i}[k]\right|.$ (8)
@@
Here index $k$ corresponds to the components within the average gradient vector $\bar{G}_{i}$ [40]. This gradient-based importance guidance addresses the key limitations of previous work. Gradient attribution provides a data-driven measure of token relevance, which can adapt to the subtle semantics of human preferences and achieve fine-grained control of key tokens in the model generation process.
Then we normalize $I_{i}$ to derive the importance-guided token-wise weight $w_{t}\in[0,1]$. Let $\mathcal{T}_{t}$ be the set of tokens in $s_{t}$. To avoid division by zero, we assume that at least two tokens have distinct importance scores in $\mathcal{T}_{t}$. The normalized weight for token $i$ is:
$w_{t}=\frac{I_{i}-\min(\mathcal{I}_{t})}{\max(\mathcal{I}_{t})-\min(\mathcal{I}_{t})},$ (9)
where $\mathcal{I}_{t}=\{I_{j}|j\in\mathcal{T}_{t}\}$ is the set of importance scores for tokens. Normalize to $[0,1]$ to maintain consistent weight measurements across diverse samples. The BT-model that $y_{\text{w}}$ is favored compared with $y_{\text{l}}$ is:
@@
Here, $w_{t}^{\text{w}}$ and $w_{t}^{\text{l}}$ are the importance weights for tokens in the preferred and less-preferred tokens, $T_{\text{w}}$ and $T_{\text{l}}$ are the lengths of $y_{\text{w}}$ and $y_{\text{l}}$ respectively. Then with $r_{\phi}(s_{t},a_{t})=\beta\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}$ in DPO, we can derive the expression for BT model:
@@
Therefore, we obtain the weighted token-level DPO base loss as [20]:

### v2

@@ Abstract
Aligning Large Language Models (LLMs) with human preferences is crucial for safe and effective AI interactions. While popular methods like Direct Preference Optimization (DPO) have simplified alignment, they remain sensitive to data noise and overlook the differential importance of individual tokens. Existing token-level approaches often rely on probability prediction or simplistic weighting schemes to obtain token importance, which still cannot fully address these issues. To solve this problem, we propose the Token-Importance Guided Direct Preference Optimization (TI-DPO), a framework that achieves fine-grained semantic control through two synergistic innovations. First, we propose a novel hybrid weighting mechanism that combines gradient attribution with a Gaussian prior, ensuring both the accuracy and robustness of token importance scores. Second, we employ a triplet loss to provide structured guidance for the optimization, explicitly guiding model outputs to approach preferred responses and diverge from non-preferred ones. Experimental results show that TI-DPO achieves higher accuracy and stronger generative diversity, providing more stable and computationally efficient solutions compared with DPO and other RLHF methods. Code and demo are available at https://github.com/gracefulning/TIDPO.
@@ 1 Introduction
Large Language Models (LLMs) have shown proficiency in Natural Language Processing (NLP) (Gao et al., 2025), logical reasoning (Xie et al., 2025a), and code generation (Xu et al., 2025), emerging as a focal point of recent research. However, as models may generate outputs inconsistent with intended purposes or ethical standards, human preference alignment aims to ensure that LLMs adhere to human values (Liu et al., 2023), producing beneficial and harmless content. Against this backdrop, Reinforcement Learning from Human Feedback (RLHF) has become a prevailing approach for achieving alignment (Hong et al., 2024; Hu et al., 2025). It leverages human-annotated preference data to train reward models and fine-tunes LLMs using Reinforcement Learning (RL) methods (Wang et al., 2023b) like Proximal Policy Optimization (PPO) (Schulman et al., 2017).
The emergence of Direct Preference Optimization (DPO) has effectively simplified the alignment process (Rafailov et al., 2023). Inspired by DPO’s implicit reward mechanism, a series of preference optimization models have been proposed in recent years, such as ORPO (Hong et al., 2024), f-DPO (Wang et al., 2023a), and CPO (Feng et al., 2025). However, both DPO and RLHF have a fundamental flaw during optimization: they optimize at the sequence level, leading to the neglect of the influence of specific tokens, which in turn destabilizes the training process due to shifts in the sampling distribution (Zhang et al., 2025).
Motivated by these challenges, researchers have proposed token-level variants of DPO, aiming to decompose preference alignment into fine-grained contributions ((Zeng et al., 2024; Xie et al., 2025a; Zhong et al., 2024)). However, achieving true fine-grained alignment requires addressing a core challenge: We not only need to accurately identify the key tokens that have a decisive impact on human preferences, but also need a subtle optimization objective to guide the model to adjust its preference (Li et al., 2025). Nevertheless, existing token-level methods fall short in dealing with this challenge for two reasons. First, their approaches to identifying key tokens often rely on biased probability proxies (Liu et al., 2024a) or overly simplified weighting schemes (Lin et al., 2024). Second, the optimization, they still inherit the binary comparison framework of DPO, simply distinguishing between “good” and “bad” samples (Meng et al., 2024). Such coarse-grained supervision signals cannot finely shape the model’s generation behavior in a continuous semantic space.
In our Token-Importance Guided Direct Preference Optimization (TI-DPO) framework, we introduce a novel hybrid weighting mechanism to accurately and robustly identify key tokens. This mechanism combines gradient attribution with a Gaussian prior, overcoming the problem of existing methods relying on biased proxies. Here, gradient attribution is a technique used to determine the contribution of each input feature (in our work, each token) to the model’s output (Ancona et al., 2017; Ballout et al., 2024). Liu et al. (2024b) offered empirical evidence that models exhibit a U-shaped attention bias, which means there is greater importance to tokens at the beginning and end of a sequence, while underweighting those in the middle. Thus, the Gaussian prior distribution here is explicitly designed to rectify this intrinsic architectural bias, ensuring that the optimization process does not neglect the semantic core of the response.
Meanwhile, we adopt a structured triplet objective based on the identified key weights to achieve fine-grained optimization by incorporating the intermediate generated outputs (Nguyen et al., 2018). This triplet structure explicitly guides the intermediate output to approach human preferences and distance from non-preferred responses, achieving fine-grained preference alignment and promoting a continuous gradient of preference learning. The mixed weights and the triplet loss complement each other and together form a complete solution for TI-DPO to achieve fine-grained alignment.
@@
  • We propose TI-DPO, a novel framework designed for achieving fine-grained alignment. This framework innovatively integrates a hybrid weighting mechanism, jointly formed by gradient attribution and a Gaussian prior, with triplet loss, significantly enhancing the robustness and stability of weight allocation.
  • Theoretically, we formalize the TI-DPO framework by providing a complete derivation of its loss function and gradient. Building on this, we prove TI-DPO achieves a tighter loss bound than DPO (Theorem 2) and the superiority of expected reward (Theorem 3). This theorem formally provides a new perspective on comprehending the superiority of TI-DPO in terms of alignment accuracy.
  • Experiment results indicate that TI-DPO surpasses existing methods in aligning LLMs with human preferences. Notably, our method achieves a leading average score of 62.3, and substantially outperforms strong baselines on key tasks such as HumanEval, TruthfulQA and IFEval with scores of 67.0, 62.0 and 75.7 respectively. Further analysis, including ablation studies and sensitivity analysis, confirms that both of our core contributions are vital to this performance.
@@ 2 Related Work
Human Preference Alignment Human preference alignment has emerged as a critical research paradigm in recent years, focusing on enabling model responses to align with human values and preferences. Early advancements mainly focused on RLHF (Ouyang et al., 2022; Bai et al., 2022) based on PPO (Schulman et al., 2017). However, these RL methods may suffer from overfitting in optimal responses. To mitigate this issue, Hu et al. (2025) introduced the Reinforce++ model, which employs batch-wise standardized rewards to prevent overfitting and enhance the prompt diversity during training. Concurrently, beyond RL approaches, Rafailov et al. (2023) introduced DPO, which obviates the need for explicit reward modeling through implicit preference learning. This implicit reward mechanism has inspired a wave of subsequent works (Cui et al., 2025), such as the SimPO algorithm proposed by Meng et al. (2024), which utilizes the sequence-averaged log probability as an implicit reward signal to streamline optimization. Notwithstanding these advancements, DPOs’ reliance on large-scale human-annotated preference datasets (Kim et al., 2025) has motivated derivative studies (Gou and Nguyen, 2024; Jiao et al., 2024) aimed at reducing data requirements. A notable example is RS-DPO (Khaki et al., 2024), which integrates rejection sampling with DPO to alleviate data scarcity. However, a more fundamental limitation pertains to the binary nature of traditional preference labels. Although the KTO method proposed by Ethayarajh et al. (2024) effectively reduces the reliance on paired preference labels in DPO, most current RLHF and related preference optimizations still mainly rely on binary comparisons between “good” and “bad” responses (Gao et al., 2024; Hong et al., 2024). Such coarse-grained supervision has obvious shortcomings: human preferences often show continuous gradient differences rather than simply “good” and “bad”. Against this backdrop, our proposed triplet optimization method can achieve fine-grained preference alignment.
@@
From Sequence-Level to Token-Level Achieving fine-grained alignment requires the model not only to distinguish the quality of the entire sequence, but also to understand and precisely control the key morphemes that constitute the semantics of the sequence. Existing sequence-level techniques often lead to a decrease in generation diversity because they ignore the importance differences among tokens (Feng et al., 2025). These limitations have spurred researchers’ research into step (Xie et al., 2024) or token-level (Rafailov et al., 2024) alignment mechanisms, seeking to address the granularity mismatch between coarse-grained sequence rewards and fine-grained token contributions (Xi et al., 2024). To address the significant decline in model generation diversity, TDPO (Zeng et al., 2024) reanalyzed and optimized the entire alignment process from the token-level perspective. An additional limitation of RLHF and DPO lies in the fact that rewards are only assigned to the final token, with all other tokens receiving no learning rewards (Zhong et al., 2024). Meanwhile, Xie et al. (2025a) proposed a correlation between the frequency of specific tokens and model performance, which inspires us to consider reassigning token weights. In a related vein, Liu et al. (2024a) estimates token importance weights using prediction probability differences. Nevertheless, this probabilistic weighting scheme is prone to bias when contrastive models produce inconsistent outputs or fail to capture subtle semantic nuances of human preferences. In contrast, our approach employs a hybrid strategy that combines causal gradient attribution with a stabilizing Gaussian prior to estimate importance (Ballout et al., 2024). By focusing on actual gradient impacts, our method enhances alignment precision over probabilistic proxies, while the prior distribution ensures robustness against noisy gradient signals.
@@ 3.1 Human Preference Alignment
Here $\sigma(\cdot)$ is the sigmoid activation function. The reward model evaluates the LLMs’ responses by applying the BT model for ranking losses (Ouyang et al., 2022):
@@
The partition function $Z(x)$ serves to normalize the policy’s probability distribution (Rafailov et al., 2023). The parameter $\beta$ regulates the extent of divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$. DPO rearranges this equation to express the reward as $r_{\phi}(x,y)=\beta\log\frac{\pi_{\theta}^{*}(y|x)}{\pi_{\text{ref}}(y|x)}-\log Z(x)$. Let the input prompt be represented as $x=[x_{1},x_{2},\ldots,x_{m}]$ and the first t1t-1 tokens generated by the model be denoted as $y^{<t}=[y^{1},y^{2},\ldots,y^{t-1}]$. Let $T_{\text{w}}$ and $T_{\text{l}}$ denote the number of preferred tokens and less preferred tokens, respectively. The token-level DPO optimization objective is given by
@@ 3.2 Triplet Loss
Here $z]_{+}$ denotes the rectified linear unit function, ensuring that it is set to zero if negative. The features extracted from the three inputs are represented by the terms $f\left({x}_{i}\right)$, $f\left({x}_{j}\right)$, and $f\left({x}_{k}\right)$.
@@ 4 Methodology
Driven by the challenges of unstable training and distribution shift in traditional RL alignment methods, we propose the TI-DPO framework. Our key innovation lies in a novel hybrid weighting strategy and a triplet loss that provides a structured optimization objective.
@@ 4.1 Token-Level MDP for LLM Preference Alignment
To address the challenges of the sequential and auto-regressive nature of text generation, a token-level Markov Decision Process (MDP) is introduced, which incorporates the notion of token significance to improve the alignment of each token selection with human preferences. This concept is defined through a tuple denoted as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho_{0})$. $\mathcal{S}$ and $\mathcal{A}$ are the state space and action space, respectively. $\mathcal{P}$ is a deterministic transition model among tokens. Here $r$ stands reward model associated with each token, and $\rho_{0}$ indicates the initial state distribution. The initial state is $s_{0}=[x]$, which is simply the input prompt. At each step $t$ of the generation process, the state $s_{t}=[x,y^{<t}]\in\mathcal{S}$ consists of input prompt $x$, where $t$ is the count of token, and t1t-1 generated tokens $y^{<t}=[y^{1},y^{2},\ldots,y^{t-1}]$. At each time step $t$, the action $a_{t}=y^{t}$ corresponds to the selection of subsequent tokens for generation.
4.2 Hybrid Weighting Mechanism for Token Importance
@@
Building on the token-level MDP framework, we formalize the calculation of importance weights $w_{t}$. Inspired by the attribution-based rationale extraction from Ballout et al. (2024), our approach quantifies token importance through gradient sensitivity analysis, ensuring that critical tokens in human-preferred responses drive the policy optimization process.
However, while gradient attribution provides a precise, data-driven signal, it can be susceptible to noise. Some studies have pointed out that imposing additional constraints on the attention or importance distribution can help the model focus on key information (Zhang et al., 2018; Guo et al., 2019). Furthermore, a recent study (Liu et al., 2024b) offered empirical evidence that models exhibit a U-shaped attention bias. This means there is greater importance to tokens at the beginning and end of a sequence, while underweighting those in the middle. The Gaussian prior distribution, which peaks at the center, is designed to counteract the architectural “Lost-in-the-Middle” bias inherent in LLMs, ensuring the optimization does not neglect the semantic core of the response. The Gaussian prior also prevents the model from overfitting to noisy gradient signals and provides a stable baseline, ensuring the weight distribution remains well-behaved throughout training.
A token that significantly impacts the reward is deemed critical, whether it has a positive effect on the preferred response or a negative effect on the non-preferred response. For a given sequence of tokens $y=[y_{1},y_{2},\dots,y_{T-1}]$, we first obtain its embedding sequence $E=[e_{1},e_{2},\dots,e_{T-1}]$, where $e_{i}$ is the embedding vector for token $y_{i}$. We then perform a forward pass to get the logits for the final token, $L_{T-1}\in\mathbb{R}^{V}$, where $V$ is the vocabulary size. The target scalar value for our gradient calculation, $\mathcal{L}_{\text{target}}$, is the maximum logit value at this final step, which represents the model’s most confident prediction for the next token $y_{T}$:
$\mathcal{L}_{\text{target}}=\max(L_{T-1}).$ (5)
Next, we compute the gradient of this target logit with respect to each token’s embedding $e_{i}$ in the sequence. This gradient, $\nabla_{e_{i}}\mathcal{L}_{\text{target}}$, captures the direct influence of token $i$ on the final prediction. To obtain a scalar importance score $I_{i}$ from the gradient vector, similar to previous work (Ballout et al., 2024), we compute its $L_{1}$ norm:
[…]
Here, $k$ indexes the components of the gradient vector. This score, $I_{i}$, represents the raw, data-driven importance of token $i$.
Finally, to ensure training stability and robustness against noise in gradient estimates, we post-process these raw scores to derive the final weights $w_{t}$. As implemented in our code, this involves a mixed strategy. First, the raw scores $\mathcal{I}=\{I_{1},\dots,I_{T-1}\}$ are normalized by their sum to form a distribution $\mathcal{I}_{\text{norm}}$. Second, we define a Gaussian-shaped prior distribution $\mathcal{P}_{\text{prior}}$ centered on the sequence, which assigns higher baseline importance to tokens in the middle. (Detailed settings for $\mathcal{P}_{\text{prior}}$ can be found in Appendix B.6.) The final weight vector $W$ is a convex combination of these two distributions, controlled by a hyperparameter $\lambda\in[0,1]$:
$W=\lambda\cdot\mathcal{I}_{\text{norm}}+(1-\lambda)\cdot\mathcal{P}_{\text{prior}}.$ (7)
This weighting scheme is applied independently to both the preferred $y_{\text{w}}$ and non-preferred $y_{\text{l}}$ sequences to obtain their respective token-level weights, denoted as $w_{t}^{\text{w}}$ and $w_{t}^{\text{l}}$.
@@
The gradient-based importance guidance method provides a data-driven measure of token relevance, which can adapt to the subtle semantics of human preferences and achieve fine-grained control over key tokens during the model generation process. These weights then modulate the implicit reward signal at each token step, effectively focusing the DPO objective on the most critical tokens. The resulting preference probability under BT model is:
@@
Here, $T_{\text{w}}$ and $T_{\text{l}}$ are the lengths of $y_{\text{w}}$ and $y_{\text{l}}$ respectively.
Then, with $r_{\phi}(s_{t},a_{t})=\beta\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}$ in DPO, we can derive the expression for BT model:
@@
Therefore, we obtain the weighted token-level DPO base loss as:
@@
4.3 Triple Loss Implementation
The practical implementation of the triplet loss is integrated seamlessly within the main training loop to provide structured guidance for the policy model.
First, for each data batch comprising $(x,y_{\text{w}},y_{\text{l}})$, the process begins by generating an anchor response $y$: Using the preferred response $y_{\text{w}}$ as the starting point of the context, the response dynamically generated by the policy model $\pi_{\theta}$ is the anchor $y$. It represents an intermediate state in the model’s generation space.
Next, by mapping each of the three responses to a point in a continuous preference space, we calculate the distances between these three responses $y,y_{\text{w}}\text{ and }y_{\text{l}}$, which represent the preference for the newly created anchor $y$.
Finally, according to the definition in Eq.(4), the triplet loss in our work is calculated with these distances, penalizing the model if the anchor is not closer to the positive response than to the negative one by a predefined margin:
$\mathcal{L}_{\text{triplet}}=\mathbb{E}_{\left(x,y_{\text{w}},y_{\text{l}}\right)\sim\mathcal{D}}\Big[\max(0,$ (12)
$\underbrace{\sum_{t=1}^{T_{\text{w}}}\Big\|\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}\!\!-\!\!\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}\Big\|^{2}_{2}}_{\text{Align }y\text{ with }y_{\text{w}}}\!\!-\!\!\underbrace{\sum_{t=1}^{T_{\text{l}}}\Big\|\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}\!\!-\!\!\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}\Big\|^{2}_{2}}_{\text{Push }y\text{ away }y_{\text{l}}}\!\!+\!\alpha)\Big]_{+}.$